# Representations of the Flat Space Wavefunction

### v1

@@ Abstract.
Supported in part by NSF grant DMS-2100288 and by Simons Foundation Collaboration Grant for Mathematicians #580839 and SFI-MPS-TSM-00012928
@@ 1. Introduction
Throughout the cosmological correlator literature, the representations mentioned above have been used implicitly. In this paper, we give proofs that $\psi_{G}$ indeed has the above canonical form representation, as well as two other representations in terms of certain collections of connected subgraphs (“tubings”), which we interpret as realizations of the bulk and boundary representations. In particular, our bulk representation is an amended version of [7, Conjecture 3.5] and equivalent to the formula proposed in [9, Equation 4.1]. Along the way, we give a regular unimodular triangulation of the dual cosmological polytope.
@@ 5. Canonical Form Representation
The toric ideal of $P^{\lor}$ is [^4] More precisely, $I_{P^{\lor}}$ is the toric ideal defined by the coefficient vectors $(f_{i1},\dots,f_{id})\in\mathbb{Z}^{d}$, where $F_{i}=f_{i1}x_{1}+\dots+f_{id}x_{d}$.
@@
Proposition 5.5 ([17, Corollary 8.4 and Theorem 8.8]).
@@
and it will be convenient to use the reverse lexicographic [^5] Note we mean reverse lexicographic, not degree reverse lexicographic. order << induced by any total order on $\{t_{T}\}$ satisfying
@@
To show that $\mathcal{G}$ is a Gröbner basis of $I_{G}$, it suffices to show that for any primitive [^7] A binomial $\mathbf{t}^{\mathbf{u}}-\mathbf{t}^{\mathbf{v}}$ in a toric ideal $I$ is primitive if there are no other binomials $\mathbf{t}^{\mathbf{r}}-\mathbf{t}^{\mathbf{s}}\in I$ with $\mathbf{t}^{\mathbf{r}}\mid\mathbf{t}^{\mathbf{u}}$ and $\mathbf{t}^{\mathbf{s}}\mid\mathbf{t}^{\mathbf{v}}$. The “suffices to show” follows from the fact that the set of all primitive binomials is always a Gröbner basis [17, Proposition 4.11]. binomial $\prod_{i}t_{S_{i}}-\prod_{j}t_{T_{j}}$ (with leading monomial $\prod_{i}t_{S_{i}}$) in $I_{G}$, the collection $\{S_{i}\}$ contains a pair of overlapping tubes. Indeed: consider the smallest monomial among $\{t_{T_{j}}\}$, say $t_{T_{0}}$, and fix a vertex $v_{0}\in T_{0}$. The equality

### v2

@@ 1. Introduction
Throughout the cosmological correlator literature, the representations mentioned above have been used implicitly. In this paper, we give proofs that $\psi_{G}$ indeed has the above canonical form representation, as well as two other representations in terms of certain collections of connected subgraphs (“tubings”), which we interpret as realizations of the bulk and boundary representations. In particular, our bulk representation is an amended version of [7, Conjecture 3.5] and equivalent to the formula proposed in [9, Equation 4.1]. Along the way, we describe a natural triangulation of the dual cosmological polytope (for triangulations of the cosmological polytope, see [13]) using the language of tubings.
@@ 5. Canonical Form Representation
The toric ideal of $P^{\lor}$ is
@@
To be precise, $I_{P^{\lor}}$ is the toric ideal defined by the coefficient vectors $\mathbf{f}_{i}\in\mathbb{Z}^{d}$ of $F_{i}$ (i.e. $F_{i}=f_{i1}x_{1}+\dots+f_{id}x_{d}$). Moreover, there is a natural multigrading on $k[t_{1},\dots,t_{m}]$ defined by $\deg(t_{i})=\mathbf{f}_{i}$, which we will make use of.
Proposition 5.5 ([18, Corollary 8.4 and Theorem 8.8]).
@@
and it will be convenient to use the reverse lexicographic [^4] Here we must use the natural multigrading on $k[t_{T}:T\in\operatorname{Tub}(G)]$, not the standard grading. order << induced by any total order on $\{t_{T}\}$ satisfying
@@
To show that $\mathcal{G}$ is a Gröbner basis of $I_{G}$, it suffices to show that for any primitive [^6] A binomial $\mathbf{t}^{\mathbf{u}}-\mathbf{t}^{\mathbf{v}}$ in a toric ideal $I$ is primitive if there are no other binomials $\mathbf{t}^{\mathbf{r}}-\mathbf{t}^{\mathbf{s}}\in I$ with $\mathbf{t}^{\mathbf{r}}\mid\mathbf{t}^{\mathbf{u}}$ and $\mathbf{t}^{\mathbf{s}}\mid\mathbf{t}^{\mathbf{v}}$. The “suffices to show” follows from the fact that the set of all primitive binomials is always a Gröbner basis [18, Proposition 4.11]. binomial $\prod_{i}t_{S_{i}}-\prod_{j}t_{T_{j}}$ (with leading monomial $\prod_{i}t_{S_{i}}$) in $I_{G}$, the collection $\{S_{i}\}$ contains a pair of overlapping tubes. Indeed: consider the smallest monomial among $\{t_{T_{j}}\}$, say $t_{T_{0}}$, and fix a vertex $v_{0}\in T_{0}$. The equality
@@
Acknowledgements
Data Availability
No data was generated or analyzed in this work.
Declarations
This work was supported in part by NSF grant DMS-2100288 and by Simons Foundation Collaboration Grant for Mathematicians #580839 and SFI-MPS-TSM-00012928. The author has no competing interests to declare.